OpenStudy (anonymous):

Is f(x) = x^2-4x-3 a function? If so, what is its domain and range? Also how do I know it's a function and its domain and range? Also how do I find x such that f(x)=2?

OpenStudy (anonymous):

if f(x)=2 x^2-4x-3=2 x^2-4x-5=0 now solve the quadratic???

OpenStudy (ricardovg):

f(x) is a function, its domain is real numbers and has a min in x=2 then is range is [-7.inf].

OpenStudy (ricardovg):

range: from f(2)=-7 to positive infinite.

OpenStudy (anonymous):

How is it a function?

OpenStudy (anonymous):

ohhh, it's a quadratic function.

OpenStudy (ricardovg):

It is a function because for each real number in the domain, you have only one value in its range. f(x) has just one value for x.

OpenStudy (anonymous):

ohh, okay. But how do I determine the domain and range?

OpenStudy (anonymous):

the domain is x belong to real numbers and range is y belongs to real number such that y must be less than or equal to -7?

OpenStudy (anonymous):

\[\left\{ x \in R \right\}\] \[\left\{ y \in R | x \le -7 \right\}\]